how to find the length and the width of the rectangle if the diagonal is 7

Answer by harpazo(655)

Use the Pythagorean Theorem.
Let x represent the length and width. If x represents the length and width, this means that your shape is actually a square NOT a rectangle. Why? Because a square has EQUAL MEASURE on all four sides.
Let d represent the diagonal.

x^2 + x^2 = d^2

2x^2 = 7^2

2x^2 = 49

x^2 = 49/2

sqrt{x^2} = sqrt{49/2}

x = 7/sqrt{2}

You may rationalize the denominator. If you do, the answer for x is
7(sqrt{2})/2. This is the measure of the length and width, which means this figure is actually a square not a rectangle.
How do we read 7(sqrt{2})/2?
The length and width is 7 times the square root of 2 divided by 2.